please provide solutions!Nittany Company uses a periodic inventory system. At the end of the annual accounting period, December 31 of the current yearHamilton Company uses a periodic inventory system. At the end of the annual accounting period, December 31 of the current yea
Answer #1

a) Calculate following

Unit Unit Cost Total Cost
Beginning inventory 2000 5 10000
Purchase 1 5000 6 30000
Purchase 2 3000 8 24000
Total 10000 64000

Calculate following

FIFO LIFO Average cost
Ending inventory (24000+6000) = 30000 (10000+2000*6) = 22000 64000/10000*4000 = 25600
Cost of goods sold 34000 42000 64000-25600 = 38400

b) Calculate following

Unit Unit Cost Total Cost
Beginning inventory 2000 5 10000
Purchase 1 6000 4 24000
Purchase 2 4000 2 8000
Total 12000 42000

Calculate following

FIFO LIFO Average cost
Ending inventory 3000*2 = 6000 10000+4000 = 14000 10500
Cost of goods sold 36000 28000 31500